Q8

For th below two machines and based on CC analysis which machine we should select? MARR=10%


Machine AMachine B
First cost, $              21,439          124,845
Annual cost, $/year                 13,160               7,237
Salvage value, $                 5,419 - 
Life, years 3infinite


Answer the below question:

B- the CC for machine B= 
Answer #1

SOLUTION :


r = MARR = 10% = 0.1

=> (1 + r) = 1.1


PV of costs of Machine A :


= 21439 + 13160/ 1.1 + 13160/1.1^2 + 13160/1.1^3 - 5419/1.1^3 

= 50094.60 ($) 


PV of costs of Machine B :


= 124845 + 7237 / 0.1

= 197215 ($) 


Therefore, as per CC, Machine A has lower cost. So Machine A should be selected.

(ANSWER)

answered by: Tulsiram Garg